# The Flavor Composition of Supernova Neutrinos

### v2

@@ Abstract
We show that standard matter effects in the outer layers of core-collapse supernovae significantly constrain the flavor composition of the neutrino flux, even with the enormous uncertainties originating from self-induced flavor conversions in the supernova core. Assuming an arbitrary initial flavor composition emerging from the supernova core, we find that, for the normal mass ordering, the fraction of neutrinos with electron flavor reaching the Earth, denoted as $f_{\nu_{e}}$, is constrained to be less than $0.5$ for all energies throughout the emission phase, whereas, for inverted mass ordering, we anticipate neutrinos arriving in near flavor equipartition ($f_{\nu_{e}}\approx 1/3$). In case adiabaticity is violated in the region of standard matter effects (e. g. due to shock waves), the result is flavor equipartition for both mass orderings. These results could be tested by future observations and may provide valuable insights into the properties of neutrino fluxes emerging from the supernova core.
@@
Introduction.— During a core-collapse supernova (SN) event, the immense gravitational binding energy of the progenitor star is predominantly liberated through a rapid emission of neutrinos, typically lasting around 10 seconds Colgate:1966ax ; Arnett:1966 . After being generated in the SN core, these neutrinos traverse the mantle and envelope, where they are thought to be pivotal in driving the explosion Wilson:1985 ; Bethe:1985sox .
@@
Figure 1: Allowed regions for the flavor composition of supernova neutrinos arriving at the Earth for the cases of normal mass ordering (left) and inverted mass ordering (right). The colors represent three different scenarios. Red: The flux emerging from collective effects is composed of flavor eigenstates in an arbitrary flavor configuration (this can be approximated by an ensemble of flavor eigenstates; see text for details), and propagation in the outer layers is adiabatic, $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. Maroon: Flux emerging from collective effects in an arbitrary flavor configuration, but non-adiabatic propagation through the $H$-resonance (presence of shock wave), $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\approx 1/3$. Stripes show overlaps between red and maroon regions. Pink: Flux of matter eigenstates in arbitrary amounts emerging from collective effects, $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\lesssim 0.7$. In this figure, we consider the $3\sigma$ range of variability only for the parameters $\theta_{12}$ and $\theta_{13}$, which affect the fraction of electron flavor $f_{\nu_{e}}$.
@@
In 1987, the detection of a few tens of SN neutrinos coming from the Large Magellanic Cloud confirmed the basic features of the core-collapse models, such as the mean energy and the duration of the neutrino burst Kamiokande-II:1987idp ; Bionta:1987qt ; Alekseev:1988gp (see Fiorillo:2023frv for a recent analysis). Nevertheless, the intricacies of the SN physics remain a challenge, largely due to the uncertainties surrounding neutrino flavor transformations in the core Tamborra:2020cul ; Ehring:2023lcd ; Ehring:2023abs ; Nagakura:2023mhr . In such a neutrino-dense environment, the coherent forward scattering of neutrinos onto each other gives rise to collective effects that drive a nonlinear flavor evolution Pantaleone:1992eq ; Pantaleone:1992xh ; Duan:2010bg ; Mirizzi:2015eza ; Chakraborty:2016yeg . These self-induced conversions are referred to as “slow” or “fast” depending on the characteristic spatial scale in which they manifest. Slow self-induced conversions develop over a length scale governed by the vacuum oscillation frequency, which is on the order of a few kilometers for typical SN neutrino energies ($\sim 10$ MeV) Duan:2010bg ; Mirizzi:2015eza . These conversions take place in an intermediate zone situated $\sim 100$ km away from the neutrino decoupling region, but before the standard matter-induced resonant neutrino conversion occurs Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . Conversely, fast self-induced conversions have characteristic scales as small as centimeters and could significantly influence the flavor evolution, energy spectrum, and angular distribution of each neutrino species near the decoupling region Sawyer:2005jk ; Sawyer:2008zs ; Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx . In addition, it was recently found that coherent flavor conversion can also be induced by incoherent collisions in the SN environment Johns:2021qby . The lack of a full solution to the quantum kinetic equations governing this collective behavior leads to unpredictable consequences for the SN dynamics Ehring:2023abs ; Nagakura:2023mhr and, ultimately, for the neutrino signal detected at the Earth Abbar:2024nhz .
@@
Amidst this conundrum, it is natural to wonder if a meaningful improvement in our understanding of the core-collapse phenomenon and neutrino propagation will be possible, even after the observation of the next galactic SN. In this Letter, we show that many features of the neutrino flux coming to the Earth can be understood by adopting a completely agnostic approach to the flavor composition emerging from the SN core, where collective effects dominate. By starting with an arbitrary flavor configuration before the onset of standard matter effects, we observe that matter effects alone impose significant constraints on the flavor composition of SN neutrinos. Assuming normal ordering (NO) of neutrino masses, we find that the fraction of electron neutrinos arriving at the Earth, denoted by $f_{\nu_{e}}^{NO}$, remains consistently below $0.5$ for all energies during the emission phase, and gradually converges to near flavor equipartition ($f_{\nu_{e}}^{NO}\approx 1/3$) [^1] In this Letter, “flavor equipartition” simply means $f_{\nu_{e}}=1/3$, since discrimination between $\nu_{\mu}$ and $\nu_{\tau}$ signals in current and next-generation detectors is challenging Akhmedov:2002zj . during the passage of shock waves. In the inverted ordering (IO) scenario, we predict flavor equipartition ($f_{\nu_{e}}^{IO}\approx 1/3$) as a generic outcome, which remains robust throughout the entire neutrino burst, even in the presence of shock waves.
@@
$\mathcal{H}_{\nu}=\frac{1}{2E}\,U\left(\begin{array}[]{ccc}0&0&0\\ 0&\Delta m_{21}^{2}&0\\ 0&0&\Delta m_{31}^{2}\end{array}\right)U^{\dagger}+\left(\begin{array}[]{ccc}V_{e}&0&0\\ 0&0&0\\ 0&0&0\end{array}\right)~{}.$ (1)
@@
$i\frac{d\rho_{\nu}}{dt}=[\mathcal{H}_{\nu},\rho_{\nu}]~{},$ (2)
where $\rho_{\nu}$ is the density matrix describing the neutrino flavor state. Observe that Eq. 2 is a simplified version of the complete expression that takes into account neutrino-neutrino interactions, advection, and collisions Sigl:1993ctk ; Cardall:2007zw ; Fiorillo:2024wej ; Fiorillo:2024fnl .
As neutrinos propagate towards vacuum and the matter density ($\rho_{m}$) gradually decreases, they encounter two resonance layers where effective flavor conversions can occur Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . These so-called $H$- and $L$-resonances occur when the following conditions are satisfied:
$\sqrt{2}G_{F}\frac{Y_{e}}{m_{N}}\rho_{m}^{\rm res}=\frac{\Delta m^{2}_{i1}\cos 2\theta_{1i}}{2E}~{},$ (3)
@@
$\gamma=\frac{\Delta m^{2}_{i1}\sin^{2}2\theta_{1i}}{2E\cos 2\theta_{1i}}\left(\frac{1}{Y_{e}\rho_{m}^{\rm res}}\left|\frac{d(Y_{e}\rho_{m})}{dr}\right|_{\rm res}\right)^{-1}~{},$ (4)
@@
Neutrino state emerging from self-induced conversions.— Assuming an emission of $10^{58}$ neutrinos over the $10$-second duration of a galactic SN burst SajjadAthar:2021prg , at a distance of 10 kpc, we estimate the flux reaching the Earth to be $N\sim 10^{15}$ neutrinos per square meter per second. Before traveling to Earth, these neutrinos reach the boundary between the region of self-induced conversions and the region of standard matter effects, where any superposition of electron neutrinos with other flavors, initially formed by self-induced conversions, becomes negligible. Therefore, at this point, the density matrix for the ensemble of $N$ neutrinos in the flavor basis is:
@@
$\rho_{\nu}^{F}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{2}&\beta_{k}\gamma_{k}e^{-i\phi_{k}}\\ 0&\beta_{k}\gamma_{k}e^{i\phi_{k}}&\gamma_{k}^{2}\end{array}\right)~{},$ (5)
@@
where $\alpha_{k}^{2}$ represents the probability of the $k$-th neutrino to interact as $\nu_{e}$, $\beta_{k}^{2}$ as $\nu_{\mu}$, and $\gamma_{k}^{2}$ as $\nu_{\tau}$, with the condition that $\alpha_{k}^{2}+\beta_{k}^{2}+\gamma_{k}^{2}=1$. The off-diagonal terms in Eq. 5 represent the interference between the $\nu_{\mu}$ and $\nu_{\tau}$ states in matter. Here, we propose that, since the neutrinos reaching the Earth may have been produced at completely different locations in the SN core, and have followed a wide variety of trajectories and evolutions before reaching the region of standard matter effects, the sum of their interference terms will destructively interfere and vanish. Based on this assumption, we can approximate the general state in Eq. 5 by an ensemble of flavor eigenstates with arbitrary proportions (i.e., $\rho_{\nu}^{F}$ is approximately diagonal) and narrow down the range of possible values for $f_{\nu_{e}}$ in the following sections.
In the Supplemental Material (SM), we show that even if destructive interference occurs in a basis other than the flavor basis, the off-diagonal terms in the flavor basis are suppressed. In this case, however, incomplete destructive interference in Eq. 5 is possible, which we treat as theoretical uncertainties in our results in upcoming sections (see SM for details).
@@
Results: Flavor ratios at the Earth.— In the following discussion, we show that under the assumption that interference terms are suppressed in Eq. 5 and that neutrino propagation is adiabatic [^3] Adiabaticity means that the relative proportions between the vacuum eigenstates that reach the Earth are identical to those of matter eigenstates propagating in the SN outer layers ($\nu_{1}^{m}\rightarrow\nu_{1}$, $\nu_{2}^{m}\rightarrow\nu_{2}$, and $\nu_{3}^{m}\rightarrow\nu_{3}$)., the $\nu_{e}$ fraction at Earth is given by $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. These results correspond to the red-shaded regions in Fig. 1.
@@
To better understand the emergence of these constraints, we characterize the evolution within the zone of standard matter effects in terms of flavor ratios $(f_{\nu_{e}},f_{\nu_{\mu}},f_{\nu_{\tau}})$. Assuming NO and adiabatic propagation, it follows that an initially produced $\nu_{e}$ arrives at the Earth as $\nu_{3}$ Dev:2023znd :
$(1,0,0)_{SN}\rightarrow\left(\left|U_{e3}\right|^{2},\left|U_{\mu 3}\right|^{2},\left|U_{\tau 3}\right|^{2}\right)_{\oplus}~{}.$ (6)
@@
Conversely, either of the non-electron flavors $\nu_{\mu}$ or $\nu_{\tau}$ arrive at the Earth approximately as equal mixtures of $\nu_{1}$ and $\nu_{2}$. Therefore,
@@
$\frac{1}{2}\left(\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2},\left|U_{\mu 1}\right|^{2}+\left|U_{\mu 2}\right|^{2},\left|U_{\tau 1}\right|^{2}+\left|U_{\tau 2}\right|^{2}\right)_{\oplus}~{}.$ (7)
For any initial combination $(a,b,c)_{SN}$ [^4] In the context of Eq. 5, we have $a=\frac{1}{N}\sum_{k}\alpha_{k}^{2}$, $b=\frac{1}{N}\sum_{k}\beta_{k}^{2}$, and $c=\frac{1}{N}\sum_{k}\gamma_{k}^{2}$., we obtain on Earth the $\nu_{e}$ fraction
$f_{\nu_{e}}^{NO}=a\left|U_{e3}\right|^{2}+b\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}+c\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}~{}.$ (8)
Because of the unitarity of the PMNS matrix and $a+b+c=1$, Eq. 8 simplifies to
$f_{\nu_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e3}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e3}\right|^{2}-1\right)~{},$ (9)
@@
$f_{\nu_{e}}^{NO}\approx\frac{(1-a)}{2}\lesssim 0.5~{}.$ (10)
The theoretical uncertainty on this upper limit is $\sigma_{\nu_{e}}^{NO}\in[-0.08,0.06]$ (see SM for details). In the IO scenario, the argument mirrors the one between Eq. 6 and Eq. 9, with the interchange of the PMNS matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ for any flavor $\alpha$. Thus, by replacing $U_{e3}$ with $U_{e2}$ in Eq. 9, we have
$f_{\nu_{e}}^{IO}=\frac{1}{2}\left(1-\left|U_{e2}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e2}\right|^{2}-1\right)~{}.$ (11)
@@
The results in Eq. 10 and Eq. 12 are applicable to neutrinos of any energy and at any instant in time. Nonetheless, exceptions may be possible when a shock wave is present, leading to modification of Eq. 10 at certain parts of the energy spectrum.
Results in the presence of shock waves.— When adiabatic propagation cannot be guaranteed, transitions between different matter eigenstates may occur at the $H$- and $L$-resonance layers. Furthermore, if the matter profile is affected by shock waves, multiple resonances of the $H$- and $L$-types could appear Schirato:2002tg ; Tomas:2004gr . If an even number of $H$-resonances are non-adiabatic, results do not change (neglecting phase effects Dasgupta:2005wn ). However, if an odd number of the $H$-resonances are non-adiabatic, the outcome is modified in the NO scenario once transitions $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ that happen at one of the resonance layers are not reverted by another transition of the same type. It is straightforward to determine the consequence of an odd number of non-adiabatic $H$-resonances: transitions of the type $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ imply that matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ should be interchanged in Eqs. 6-9. The argument is similar to the one that leads to Eqs. 11 and Eq. 12, resulting in
$f_{\nu_{e}}^{NO}(\text{non-ad $H$})\approx\frac{1}{3}~{}.$ (13)
Because the position of the $H$-resonance is energy dependent, the portion of the energy spectrum influenced by the transient non-adiabaticity provides information about the location of the shock wave front Schirato:2002tg ; Tomas:2004gr .
@@
$f_{\nu_{e}}^{IO}(\text{non-ad $H$})\approx\frac{1}{3}~{}.$ (14)
@@
Although shock wave fronts can also penetrate the $L$-resonance layer, adiabaticity is never strongly violated there Mirizzi:2015eza . Even in cases where such violations may happen, they typically occur at later times (beyond $10$ seconds post-bounce) Tomas:2004gr , so we do not consider them here.
@@
Additional remarks.— The extreme values of the interval in Eq. 10 suggest that neutrinos would emerge from the inner regions either entirely in the electron flavor ($f_{\nu_{e}}^{NO}\approx 0$) or completely devoid of it ($f_{\nu_{e}}^{NO}\approx 0.5$). These scenarios are probably unlikely, as they would require a specific mechanism to produce neutrinos in these extreme states [^6] Exception is the early stage of emission Serpico:2011ir , during which self-induced conversions are negligible and neutrinos emerge mostly as $\nu_{e}$.. Therefore, the $f_{\nu_{e}}^{NO}$ range should often approach the one for IO (Eq. 12).
@@
The results above emerge from the maximal mixing between $\nu_{\mu}$ and $\nu_{\tau}$ in matter, which restricts two of the three mass eigenstates to reach the Earth in similar numbers (Eq. The Flavor Composition of Supernova Neutrinos). It is interesting to contrast the results above with the situation where any combination of mass eigenstates is allowed, $f_{\nu_{e}}\lesssim 0.7$ Horiuchi:2018ofe , given by the pink region in Fig. 1.
@@
Conclusion.— We have shown that the flavor composition of SN neutrinos reaching the Earth can be constrained, even when assuming complete uncertainty about the neutrino state that emerges from the innermost regions dominated by self-induced conversions. For NO, we predict that the proportion of $\nu_{e}$ is always smaller than $0.5$, whereas, for IO, we expect near flavor equipartition ($f_{\nu_{e}}\approx 1/3$).
@@
Ultimately, this Letter advocates for a phenomenological approach to comprehensively address the intricacies of neutrino flavor conversion in SN environments. Experimental challenges along the way include distinguishing between $\nu_{x}$ and $\bar{\nu}_{x}$ in neutral current scatterings Jachowicz:2004we . This approach can also be extended to accommodate non-standard scenarios, potentially unveiling novel phenomena underlying the physics of neutrino masses and mixing Jana:2022tsa ; Jana:2023ufy .

### v3

# Flavor composition of supernova neutrinos

@@ Abstract
Predicting the flavor composition of neutrinos from supernovae is a challenging task, primarily due to the high neutrino densities at their core. In such an environment, neutrino self-interactions give rise to collective effects that have dramatic yet poorly understood consequences for their flavor evolution. In this paper, however, we show that standard matter effects in the outer layers of supernovae can significantly constrain the flavor composition of the neutrino flux. We assume that, since a large number of neutrinos undergo different evolutions within the core, their state upon entering the MSW-dominated region is affected by decoherence. This assumption simplifies the problem and suggests that the fraction of neutrinos with electron flavor reaching Earth, denoted as $f_{\nu_{e}}$, is constrained to be less than $0.5$ for all energies throughout the emission phase in the case of normal mass ordering. In contrast, for inverted mass ordering, we anticipate neutrinos arriving in near flavor equipartition ($f_{\nu_{e}}\approx 1/3$). These predictions, and consequently their underlying assumptions, could be tested by future observations and may provide valuable insights into the properties of neutrino fluxes emerging from supernovae.
@@
Introduction.— During a core-collapse supernova (SN) event, the immense gravitational binding energy of the progenitor star is predominantly liberated through a rapid emission of neutrinos, typically lasting around 10 s Colgate:1966ax ; Arnett:1966 . After being generated in the SN core, these neutrinos traverse the mantle and envelope, where they are thought to be pivotal in driving the explosion Wilson:1985 ; Bethe:1985sox .
@@
Figure 1: Allowed regions for the flavor composition of supernova neutrinos arriving at Earth for the cases of normal mass ordering (left) and inverted mass ordering (right). The colors represent three different scenarios. Red: Flux emerging from collective effects in an arbitrary flavor configuration (this can be approximated by an ensemble of flavor eigenstates; see text for details), and propagation in the outer layers is adiabatic, $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. Maroon: Flux emerging from collective effects in an arbitrary flavor configuration, but nonadiabatic propagation through the $H$-resonance (presence of shock wave), $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\approx 1/3$. Stripes show overlaps between red and maroon regions. Pink: Flux of matter eigenstates in arbitrary amounts emerging from collective effects, $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\lesssim 0.7$. In this figure, we consider the $3\sigma$ range of variability only for the parameters $\theta_{12}$ and $\theta_{13}$, which affect the fraction of electron flavor $f_{\nu_{e}}$.
@@
In 1987, the detection of a few tens of SN neutrinos coming from the Large Magellanic Cloud confirmed the basic features of the core-collapse models, such as the mean energy and the duration of the neutrino burst Kamiokande-II:1987idp ; Bionta:1987qt ; Alekseev:1988gp (see Fiorillo:2023frv for a recent analysis). Nevertheless, the intricacies of the SN physics remain a challenge, largely due to the uncertainties surrounding neutrino flavor transformations in the core Tamborra:2020cul ; Ehring:2023lcd ; Ehring:2023abs ; Nagakura:2023mhr . In such a neutrino-dense environment, the coherent forward scattering of neutrinos onto each other gives rise to collective effects that drive a nonlinear flavor evolution Pantaleone:1992eq ; Pantaleone:1992xh ; Duan:2010bg ; Mirizzi:2015eza ; Chakraborty:2016yeg . These self-induced conversions are referred to as “slow” or “fast” depending on the characteristic spatial scale in which they manifest. Slow self-induced conversions develop over a length scale governed by the vacuum oscillation frequency, which is on the order of a few kilometers for typical SN neutrino energies ($\sim 10$ MeV) Duan:2010bg ; Mirizzi:2015eza . These conversions take place in an intermediate zone situated $\sim 100$ km away from the neutrino decoupling region, but before the standard matter-induced resonant neutrino conversion occurs Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . Conversely, fast self-induced conversions have characteristic scales as small as centimeters and could significantly influence the flavor evolution, energy spectrum, and angular distribution of each neutrino species near the decoupling region Sawyer:2005jk ; Sawyer:2008zs ; Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx . In addition, it was recently found that coherent flavor conversion can also be induced by incoherent collisions in the SN environment Johns:2021qby . The lack of a full solution to the quantum kinetic equations governing this collective behavior leads to unpredictable consequences for the SN dynamics Ehring:2023abs ; Nagakura:2023mhr and, ultimately, for the neutrino signal detected at Earth Abbar:2024nhz .
@@
Amidst this conundrum, it is natural to wonder if a meaningful improvement in our understanding of the core-collapse phenomenon and neutrino propagation will be possible, even after the observation of the next galactic SN. In this paper, we show that many features of the neutrino flux coming to Earth can be understood by making a simple assumption regarding the state of the neutrino ensemble as they emerge from the SN core. Namely, due to the large number of neutrinos interfering randomly, we expect coherence to be lost, rendering the density matrix for the ensemble diagonal in some basis (which need not be the flavor basis or any specific basis a priori) as they enter the MSW-dominated region. We then show that, for most of these bases, matter effects alone impose significant constraints on the flavor composition of supernova neutrinos. Assuming normal ordering (NO) of neutrino masses, the fraction of electron neutrinos arriving at Earth, denoted by $f_{\nu_{e}}^{NO}$, is found to remain consistently below $0.5$ for all energies during the emission phase, and gradually converges to near flavor equipartition ($f_{\nu_{e}}^{NO}\approx 1/3$) [^1] For our purpose, “flavor equipartition” simply means $f_{\nu_{e}}=1/3$, since discrimination between $\nu_{\mu}$ and $\nu_{\tau}$ signals in current and next-generation detectors is challenging Akhmedov:2002zj . during the passage of shock waves. In the inverted ordering (IO) scenario, we predict flavor equipartition ($f_{\nu_{e}}^{IO}\approx 1/3$) as a generic outcome, which remains robust throughout the entire neutrino burst, even in the presence of shock waves.
@@
$\mathcal{H}_{\nu}=\frac{1}{2E}\,U\left(\begin{array}[]{ccc}0&0&0\\ 0&\Delta m_{21}^{2}&0\\ 0&0&\Delta m_{31}^{2}\end{array}\right)U^{\dagger}+\left(\begin{array}[]{ccc}V_{e}&0&0\\ 0&0&0\\ 0&0&0\end{array}\right).$ (1)
@@
$i\frac{d\rho_{\nu}}{dt}=[\mathcal{H}_{\nu},\rho_{\nu}],$ (2)
where $\rho_{\nu}$ is the density matrix describing the neutrino flavor state. Observe that Eq. (2) is a simplified version of the complete expression that takes into account neutrino-neutrino interactions, advection, and collisions Sigl:1993ctk ; Cardall:2007zw ; Fiorillo:2024wej ; Fiorillo:2024fnl .
As neutrinos propagate toward vacuum and the matter density ($\rho_{m}$) gradually decreases, they encounter two resonance layers where effective flavor conversions can occur Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . These so-called $H$- and $L$-resonances occur when the following conditions are satisfied:
$\sqrt{2}G_{F}\frac{Y_{e}}{m_{N}}\rho_{m}^{\rm res}=\frac{\Delta m^{2}_{i1}\cos 2\theta_{1i}}{2E},$ (3)
@@
$\gamma=\frac{\Delta m^{2}_{i1}\sin^{2}2\theta_{1i}}{2E\cos 2\theta_{1i}}\left(\frac{1}{Y_{e}\rho_{m}^{\rm res}}\left|\frac{d(Y_{e}\rho_{m})}{dr}\right|_{\rm res}\right)^{-1},$ (4)
@@
Neutrino state emerging from self-induced conversions.— Assuming an emission of $10^{58}$ neutrinos over the $10$ s duration of a galactic SN burst SajjadAthar:2021prg , at a distance of 10 kpc, we estimate the flux reaching Earth to be $N\sim 10^{15}$ neutrinos per square meter per second. Before traveling to Earth, these neutrinos reach the boundary between the region of self-induced conversions and the region of standard matter effects, where any superposition of electron neutrinos with other flavors, initially formed by self-induced conversions, becomes negligible. Therefore, at this point, the density matrix for the ensemble of $N$ neutrinos in the flavor basis $F$ is
@@
$\rho_{\nu}^{F}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{2}&\beta_{k}\gamma_{k}e^{-i\phi_{k}}\\ 0&\beta_{k}\gamma_{k}e^{i\phi_{k}}&\gamma_{k}^{2}\end{array}\right),$ (5)
@@
where $\alpha_{k}^{2}$ represents the probability of the $k$th neutrino to interact as $\nu_{e}$, $\beta_{k}^{2}$ as $\nu_{\mu}$, and $\gamma_{k}^{2}$ as $\nu_{\tau}$, with the condition that $\alpha_{k}^{2}+\beta_{k}^{2}+\gamma_{k}^{2}=1$. The off-diagonal terms in Eq. (5) represent the interference between the $\nu_{\mu}$ and $\nu_{\tau}$ states in matter. Here, we propose that, since the neutrinos reaching Earth may have been produced at completely different locations in the SN core, and have followed a wide variety of trajectories and evolutions before reaching the region of standard matter effects, the sum of their interference terms will destructively interfere and vanish in some basis $B$. In the appendix, we show that even if destructive interference occurs in a basis other than the flavor basis, the off-diagonal terms in the flavor basis are suppressed. [^3] There is no reason to assume that there is a preferred basis in which the complex contributions to the off-diagonal of Eq. 5 average to zero, nor that this basis remains the same at different times during the emission. Based on this assumption, we can approximate the general state in Eq. (5) by an ensemble of flavor eigenstates with arbitrary proportions (i.e., $\rho_{\nu}^{F}$ is approximately diagonal) and narrow down the range of possible values for $f_{\nu_{e}}$ in the following sections. The reader should keep in mind that, if $B\neq F$, incomplete destructive interference in Eq. (5) is still possible. We treat this as theoretical uncertainties in our results in the upcoming sections (see the appendix for details).
@@
Results: Flavor ratios at the Earth.— In the following discussion, we show that under the assumption that interference terms are suppressed in Eq. (5) and that neutrino propagation is adiabatic, the $\nu_{e}$ fraction at Earth is given by $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. These results correspond to the red-shaded regions in Fig. 1.
@@
To better understand the emergence of these constraints, we characterize the evolution within the zone of standard matter effects in terms of flavor ratios $(f_{\nu_{e}},f_{\nu_{\mu}},f_{\nu_{\tau}})$. Assuming NO and adiabatic propagation, it follows that an initially produced $\nu_{e}$ arrives at Earth as $\nu_{3}$ Dev:2023znd :
$(1,0,0)_{SN}\rightarrow\left(\left|U_{e3}\right|^{2},\left|U_{\mu 3}\right|^{2},\left|U_{\tau 3}\right|^{2}\right)_{\oplus}.$ (6)
@@
Conversely, either of the nonelectron flavors $\nu_{\mu}$ or $\nu_{\tau}$ arrive at the Earth approximately as equal mixtures of $\nu_{1}$ and $\nu_{2}$. Therefore,
@@
$\frac{1}{2}\left(\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2},\left|U_{\mu 1}\right|^{2}+\left|U_{\mu 2}\right|^{2},\left|U_{\tau 1}\right|^{2}+\left|U_{\tau 2}\right|^{2}\right)_{\oplus}.$ (7)
For any initial combination $(a,b,c)_{SN}$, [^4] In the context of Eq. (5), we have $a=\frac{1}{N}\sum_{k}\alpha_{k}^{2}$, $b=\frac{1}{N}\sum_{k}\beta_{k}^{2}$, and $c=\frac{1}{N}\sum_{k}\gamma_{k}^{2}$. we obtain on Earth the $\nu_{e}$ fraction
$f_{\nu_{e}}^{NO}=a\left|U_{e3}\right|^{2}+b\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}+c\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}.$ (8)
Because of the unitarity of the PMNS matrix and $a+b+c=1$, Eq. (8) simplifies to
$f_{\nu_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e3}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e3}\right|^{2}-1\right).$ (9)
@@
$f_{\nu_{e}}^{NO}\approx\frac{(1-a)}{2}\lesssim 0.5.$ (10)
The theoretical uncertainty on this upper limit is $\sigma_{\nu_{e}}^{NO}\in[-0.08,0.06]$ (see the appendix for details). In the IO scenario, the argument mirrors the one between Eq. (6) and (9), with the interchange of the PMNS matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ for any flavor $\alpha$. Thus, by replacing $U_{e3}$ with $U_{e2}$ in Eq. (9), we have
$f_{\nu_{e}}^{IO}=\frac{1}{2}\left(1-\left|U_{e2}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e2}\right|^{2}-1\right).$ (11)
@@
The results in Eq. (10) and (12) are applicable to neutrinos of any energy and at any instant in time. Nonetheless, exceptions may be possible when a shock wave is present, leading to modification of Eq. (10) at certain parts of the energy spectrum.
Results in the presence of shock waves.— When adiabatic propagation cannot be guaranteed, transitions between different matter eigenstates may occur at the $H$- and $L$-resonance layers. Furthermore, if the matter profile is affected by shock waves, multiple resonances of the $H$- and $L$-types could appear Schirato:2002tg ; Tomas:2004gr . If an even number of $H$-resonances are nonadiabatic, results do not change [neglecting phase effects Dasgupta:2005wn ]. However, if an odd number of the $H$-resonances are nonadiabatic, the outcome is modified in the NO scenario once transitions $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ that happen at one of the resonance layers are not reverted by another transition of the same type. It is straightforward to determine the consequence of an odd number of nonadiabatic $H$-resonances: transitions of the type $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ imply that matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ should be interchanged in Eqs. (6)-(9). The argument is similar to the one that leads to Eqs. (11) and (12), resulting in
$f_{\nu_{e}}^{NO}(\text{non-ad $H$})\approx\frac{1}{3}.$ (13)
Because the position of the $H$-resonance is energy dependent, the portion of the energy spectrum influenced by the transient nonadiabaticity provides information about the location of the shock wave front Schirato:2002tg ; Tomas:2004gr .
@@
$f_{\nu_{e}}^{IO}(\text{non-ad $H$})\approx\frac{1}{3}.$ (14)
@@
Although shock wave fronts can also penetrate the $L$-resonance layer, adiabaticity is never strongly violated there Mirizzi:2015eza . Even in cases where such violations may happen, they typically occur at later times (beyond $10$ s post-bounce) Tomas:2004gr , so we do not consider them here.
@@
Additional remarks.— The extreme values of the interval in Eq. (10) suggest that neutrinos would emerge from the inner regions either entirely in the electron flavor ($f_{\nu_{e}}^{NO}\approx 0$) or completely devoid of it ($f_{\nu_{e}}^{NO}\approx 0.5$). These scenarios are probably unlikely, as they would require a specific mechanism to produce neutrinos in these extreme states. [^6] Exception is the early stage of emission Serpico:2011ir , during which self-induced conversions are negligible and neutrinos emerge mostly as $\nu_{e}$. Therefore, the $f_{\nu_{e}}^{NO}$ range should often approach the one for IO [Eq. (12)].
@@
The results above emerge from the maximal mixing between $\nu_{\mu}$ and $\nu_{\tau}$ in matter, which restricts two of the three mass eigenstates to reach Earth in similar numbers [Eq. (Flavor composition of supernova neutrinos)]. It is interesting to contrast the results above with the situation where any combination of mass eigenstates is allowed, $f_{\nu_{e}}\lesssim 0.7$ Horiuchi:2018ofe , given by the pink region in Fig. 1.
@@
Conclusion.— We have shown that the flavor composition of SN neutrinos reaching Earth can be constrained due to the distinct flavor evolution of neutrinos in the core, which leads to the decoherence of the neutrino ensemble upon reaching the MSW-dominated outer regions. For NO, we predict that the proportion of $\nu_{e}$ is always smaller than $0.5$, whereas for IO, we expect near flavor equipartition ($f_{\nu_{e}}\approx 1/3$).
@@
Ultimately, this article advocates for a phenomenological approach to comprehensively address the intricacies of neutrino flavor conversion in SN environments. Experimental challenges along the way include distinguishing between $\nu_{x}$ and $\bar{\nu}_{x}$ in neutral current scatterings Jachowicz:2004we . This approach can also be extended to accommodate non-standard scenarios, potentially unveiling novel phenomena underlying the physics of neutrino masses and mixing Jana:2022tsa ; Jana:2023ufy .